What is an equation of the line which is parallel to 3x+7y=4 and passes
through (5,-2) ?

Found 3 solutions by Boreal, Edwin McCravy, MathTherapy:
Answer by Boreal(15235)

Get the slope of the first line, which is 7y=-3x+4 and y=-(3/7)x+(4/7)
The slope of the line desired is -3/7
point slope formula y-y1=m(x-x1), m slope (x1, y1) point
y+2=(-3/7)*(x-5)
y=-(3/7)x+(15/7)-2
=(-3/7)x+(1/7)



Answer by Edwin McCravy(20055)

Any line parallel to to the line ax+by=c can have equation ax+by = K, where
K ≠ c

What is an equation of the line which is parallel to 3x+7y = 4 and passes
through (5,-2)
It can have the equation 3x+7y = K where K ≠ 4

To find K, simply substitute x = 5 and y = -2 in

     3x+7y = K  
3(5)+7(-3) = K
     15-21 = K
       -16 = K

Answer:  3x+7y = -16

Edwin

Answer by MathTherapy(10552)


What is an equation of the line which is parallel to 3x+7y=4 and passes
through (5,-2) ?